a 4.0 kg object is hanging from a vertical rope. the upper end of the rope is at rest. how must the person move threir hand so that the tention is equal to 44N?

a. the string is at rest

b. must be 11 m/s upward

c. must be 11m/s2 upward

d.must be 1.2 m/s2 upward

e.must be 1.2 m/s2 downward
Answer #1

Applying the newton's second law,

As tension is acting upwards and mg is acting downwards, we can write the equation as

T - mg = ma , where a is in upward direction.

44 - 4*9.8 = 4*a

a = 11 - 9.8 = 1.2 m/s2

Therefore, its acceleration must be 1.2 m/s2 upwards. i.e. correct answer should be (d).

I hope i was able to answer your question. If you need anymore explanation, please ask in comments. I would be happy to help.

Please rate the solution if you are satisfied with the solution.